Circular No. 6585

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) H. Matsuo, A. Sakamoto, N. Kuno, and N. Ukita, Nobeyama Radio

Observatory; and T. Fuse, Graduate University for Advanced Studies, report: “Millimeter-wave continuum observations of comet C/1995 O1 at 150 GHz, with the Nobeyama 45-m telescope (+ seven-element bolometer array) on Mar. 2.96 and 4.92 UT, yield detection of a compact component with a flux density of 55 ± 7 mJy. We also observed the comet on Mar. 5.92 at 100 GHz with a 4-beam SIS receiver and measured the flux density as 29 ± 4 mJy. Lower-frequency observations are highly desired for modelling the dust spectrum of the comet.”

1RXS J170849.0–400910 M. Sugizaki and F. Nagase, Institute of Space and Astronautical Sci-

ence; S. Yamauchi, Iwate University; K. Koyama, Kyoto University; K. Matsuzaki, University of Tokyo; T. Asanuma, K. Kinugasa, and K. Torii, Osaka University; and T. Kotani, Riken, communicate on behalf of the ASCA Galactic Ridge Survey Team: “During the galactic ridge survey with ASCA, we detected a relatively bright, soft source on 1996 Sept. 3.021–3.590 UT at α = 17^{h}08^{m}46^{s}, δ = −40^{o}09′27′′ (equinox 2000.0). This source is identified with the ROSAT source 1RXS J170849.0–400910 within the ASCA position determination accuracy of radius 0′.5. From timing analysis of the source, we discovered coherent pulsations with a barycentric pulse period P = 10.997 ± 0.001 s. The x-ray pulse profile shows a broad sinusoidal shape with a pulse fraction of ∼ 30 percent. The energy spectrum in the range 1–10 keV is very soft and can be fitted by a power-law model with photon index 3.5 and an absorption column density 1.8 × 10^{22} cm^{−2}. The energy spectrum can also be fitted with an optically thin thermal bremsstrahlung model with kT = 1.8 keV, although no emission line of iron is found with an upper limit of equivalent width 500 eV (either for a 6.4-, 6.6-, or 6.9-keV line; 90-percent confidence limit). The pulse-phase averaged flux in the range 1–10 keV is ∼ 1.2 × 10^{−10} cm^{−2} s^{−1} after correcting for soft x-ray absorption. From the 11-s period and the soft spectrum, we suspect this source to be the sixth member of a class of braking x-ray pulsars (Mereghetti and Stella 1995, Ap.J. 442, L17; van Paradijs et al. 1995, A.Ap. 299, L41).”

1997 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green